You wish to test the following claim (HaHa) at a significance level of α=0.01.

      Ho:μ=72.5
      Ha:μ<72.5

You believe the population is normally distributed, but you do not know the standard deviation. You obtain a sample of size n=28 with mean M=66.9 and a standard deviation of SD=10.5.

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value =

The p-value is...

  • less than (or equal to) αα
  • greater than αα



This p-value leads to a decision to...

  • reject the null
  • accept the null
  • fail to reject the null



As such, the final conclusion is that...

  • There is sufficient evidence to warrant rejection of the claim that the population mean is less than 72.5.
  • There is not sufficient evidence to warrant rejection of the claim that the population mean is less than 72.5.
  • The sample data support the claim that the population mean is less than 72.5.
  • There is not sufficient sample evidence to support the claim that the population mean is less than 72.5.
Answer #1

Test statistic t= (xbar - )/(s/√n)

t = ( 66.9- 72.5)/(10.5/√28) = -2.82

p-value for t = -2.82 and d.f = n -1 = 27

p-value = P( t < -2.82) d.f = 27

p-value = 0.0044

For a = 0.01

p-value is less than a

p-value leads the decision reject the null

Conclusion :

  • There is not sufficient evidence to warrant rejection of the claim that the population mean is less than 72.5.